makylah65:

The table shows the relationship between the number of cups of cereal and the number of cups of raisins in a cereal bar recipe: Number of cups of cereal Number of cups of raisins 6 2 9 3 12 4 Jake plotted an equivalent ratio on the coordinate grid below: A graph is shown. The title of the graph is Cereal Bars. The horizontal axis label is Cereal in cups. The horizontal axis values are 0, 30, 60, 90, 120, 150. The vertical axis label is Raisins in cups. The vertical axis values are 0, 20, 40, 60, 80, 100. A point is plotted on the ordered pair 30, 20. Which of these explains why the point Jake plotted is incorrect?

Sprinkles285:

15 cups of cereal to 5 cups of raisins = 60 cups of cereal to x cups of raisins 15/5 = 60/x -- cross multiply 15x = (60)(5) 15x = 300 x = 300/15 x = 20 so if there are 60 cups of cereal, then there are 20 cups of raisins. I think I did that right. If not then shoot me a dm.

Sprinkles285:

I hope that answers your question

Aphrodite285:

We all appreciate each other for all the kind things we do on this app!

Aphrodite285:

Thank you for helping @makylah65 !